OpenStudy (anonymous):

Simplify expression. Assume that all variables represent positive real numbers. (x^2y^6) 1/2 over (-2xy)^3

OpenStudy (anonymous):

\[\frac{1}{-8x^2}

OpenStudy (anonymous):

\[\frac{1}{-8x^2}\]

OpenStudy (anonymous):

Is it \(\frac{1}{2} (x^2y^2) \div (-2xy)^3\)?

OpenStudy (anonymous):

I meant \(y^6\)

OpenStudy (anonymous):

the 1/2 is after not before

OpenStudy (anonymous):

the half is the exponent

OpenStudy (anonymous):

Yes AnwarA it is.